# Activated LoRA: Fine-tuned LLMs for Intrinsics

### v4

@@ F.3 Training and evaluation details
Following typical LoRA SFT best practices, for both LoRA and aLoRA we used 4 training epochs, alpha of 32, dropout of 0.05, adapted the K, Q, and V modules in all layers, and searched over ranks $[6,8,16,32]$ and learning rates $[3\times 10^{-6},10^{-5},3\times 10^{-5},10^{-4},3\times 10^{-4}]$. Batch size of 8 was used, with 16-bit arithmetic precision. Hyperparameters were selected by taking the configuration that performed best on the validation set, and reported performance was computed on the test set for those selected models. All training runs were done on single H100 GPUs.
@@ Appendix G Additional details for intrinsics experiments
For the intrinsics tasks, all attention weights (keys, queries, values) were adapted in all layers, using rank 32 adapters. The learning rate and number of epochs were tuned to achieve the best validation performance (as was the case for the LoRA adapters of Danilevsky et al. [2025]). Intrinsics training tasks each used an 8 GPU H100 node.
@@ G.1 Uncertainty Quantification
Certainty score interpretation The returned percentages are calibrated in the following sense: given a set of answers assigned a certainty score of X%, approximately Xof these answers should be correct. Here “approximately" can be quantified via the expected calibration error, or ECE. Essentially what happens is teaching the adapter model what the base model knows and doesn’t know. This inherently requires generalization to questions of wildly varying difficulty (some of which may be trick questions!) and to settings not in training. Intuitively, it does this by extrapolating based on related questions it has been evaluated on in training - this is an inherently inexact process and leads to some hedging. Training pipeline First, a probe-based model was trained to produce calibrated certainty scores, using a large diverse collection of QA datasets detailed in Appendix H.1. Note that throughout, the chat template was used. The procedure for this was as follows:
@@
The above follows the procedure of Shen et al. [2024a] for freeform responses, and was applied to both the multiple choice and freeform data for consistency.
@@ G.2 Answerability determination
The input to the model is a list of conversational turns and a list of documents converted to a string using apply_chat_template function. These turns can alternate between the user and assistant roles. The last turn is from the user. The list of documents is a dictionary with text field, which contains the text of the corresponding document. To prompt the aLoRA adapter to determine answerability, a special answerability role is used to trigger this capability of the model. The role includes the keyword "answerability": <|start_of_role|>answerability<|end_of_role|> When prompted with the above input, the model generates the answerable or unanswerable output. See Danilevsky et al. [2025] for more details.
@@ H.3 Training data for Answerability Determination Intrinsic
The training data uses the publicly available Government corpus from MT-RAG Katsis et al. [2025] as the source of documents. Based on this corpus, the dataset consists of a mix of human-created and synthetically generated multi-turn conversations. It includes two types of examples: (1) Answerable queries, where the final user question can be answered based on the provided documents. These examples teach the adapter to recognize when sufficient information is present to support an answer. (2) Unanswerable queries, where the documents lack the necessary information to answer the final user query. Mixtral was used as an automatic judge to validate the answerability labels and filter out noisy samples.
@@ H.4 Training data for Jailbreak Intrinsic
The Jailbreak Intrinsic was trained on 40,000 harmful and benign samples. This is composed of several sub-sampled open source datasets (Gandalf Ignore Instructions Pfister et al. [2025], Awesome ChatGPT Prompts awe , BoolQ Clark et al. [2019], SAP Deng et al. [2023], UltraChat Ding et al. [2023], super natural instructions Wang et al. [2022b]) as well as non-public prompt datasets of harmful/benign content.

### v5

@@ F.3 Training and evaluation details
Following typical LoRA SFT best practices, for both LoRA and aLoRA we used 4 training epochs, alpha of 32, dropout of 0.05, adapted the K, Q, and V modules in all layers, and searched over ranks $[6,8,16,32]$ and learning rates $3\times 10^{-6},10^{-5},3\times 10^{-5},10^{-4},3\times 10^{-4}]$. Batch size of 8 was used, with 16-bit arithmetic precision. Hyperparameters were selected by taking the configuration that performed best on the validation set, and reported performance was computed on the test set for those selected models. All training runs were done on single H100 GPUs.
@@ Appendix G Additional details for intrinsics experiments
For the intrinsics tasks, all attention weights (keys, queries, values) were adapted in all layers, using rank 32 adapters. The learning rate and number of epochs were tuned to achieve the best validation performance (as was the case for the LoRA adapters of [5]). Intrinsics training tasks each used an 8 GPU H100 node.
@@ G.1 Uncertainty Quantification
Certainty score interpretation The returned percentages are calibrated in the following sense: given a set of answers assigned a certainty score of X%, approximately X% of these answers should be correct. Here “approximately" can be quantified via the expected calibration error, or ECE. Essentially what happens is teaching the adapter model what the base model knows and doesn’t know. This inherently requires generalization to questions of wildly varying difficulty (some of which may be trick questions!) and to settings not in training. Intuitively, it does this by extrapolating based on related questions it has been evaluated on in training - this is an inherently inexact process and leads to some hedging. Training pipeline First, a probe-based model was trained to produce calibrated certainty scores, using a large diverse collection of QA datasets detailed in Appendix H.1. Note that throughout, the chat template was used. The procedure for this was as follows:
@@
The above follows the procedure of [32] for freeform responses, and was applied to both the multiple choice and freeform data for consistency.
@@ G.2 Answerability determination
The input to the model is a list of conversational turns and a list of documents converted to a string using apply_chat_template function. These turns can alternate between the user and assistant roles. The last turn is from the user. The list of documents is a dictionary with text field, which contains the text of the corresponding document. To prompt the aLoRA adapter to determine answerability, a special answerability role is used to trigger this capability of the model. The role includes the keyword "answerability": <|start_of_role|>answerability<|end_of_role|> When prompted with the above input, the model generates the answerable or unanswerable output. See [5] for more details.
@@ H.3 Training data for Answerability Determination Intrinsic
The training data uses the publicly available Government corpus from MT-RAG [15] as the source of documents. Based on this corpus, the dataset consists of a mix of human-created and synthetically generated multi-turn conversations. It includes two types of examples: (1) Answerable queries, where the final user question can be answered based on the provided documents. These examples teach the adapter to recognize when sufficient information is present to support an answer. (2) Unanswerable queries, where the documents lack the necessary information to answer the final user query. Mixtral was used as an automatic judge to validate the answerability labels and filter out noisy samples.
@@ H.4 Training data for Jailbreak Intrinsic
The Jailbreak Intrinsic was trained on 40,000 harmful and benign samples. This is composed of several sub-sampled open source datasets (Gandalf Ignore Instructions [27], Awesome ChatGPT Prompts [1], BoolQ [3], SAP [6], UltraChat [8], super natural instructions [42]) as well as non-public prompt datasets of harmful/benign content.